OpenStudy (openstudymathhelp):

g(x)=sqrt(x)*(2x^2-4) when x=4 Derivatives. I have to find the equation of the tangent line.

OpenStudy (openstudymathhelp):

\[g(x)=\sqrt{x}(2x^2-4)\]

sam (.sam.):

Distribute it then product rule afterwards?

OpenStudy (openstudymathhelp):

I just feel like my slope is wrong. Especially for the equation

OpenStudy (openstudymathhelp):

Or do i plug in the 4 once I find the derivative to find the slope?

sam (.sam.):

Once you differentiate that that function and plugged x=4 in, you will get your slope \(\frac{dy}{dx}\)

sam (.sam.):

\[g(x)=\sqrt{x} \left(2 x^2-4\right)\] \[g'(x)=\frac{5 x^2-2}{\sqrt{x}}\]

OpenStudy (openstudymathhelp):

Then how would I find the y intercept? Do I plug the x back into the original equation?

sam (.sam.):

When x=4, \[g'(x)=\frac{5 x^2-2}{\sqrt{x}}=39\]

OpenStudy (openstudymathhelp):

Yeah I got 39 as well for the slope.

sam (.sam.):

Use x=4 on g(x) to find the 'y' value, then once you've got a coordinate, and a slope, you can create an equation \((y-y_1)=m(x-x_1)\)

OpenStudy (openstudymathhelp):

Thank you so much. This makes so much more sense now!

sam (.sam.):

Cheers

OpenStudy (phi):

Here is a graph of the tangent line